# MemeBlip2: A Novel Light Weight Multimodal System to Detect Harmful Memes thanks: :

### v1

@@ Abstract
Memes often merge visuals with brief text to share humor or opinions, yet some memes contain harmful messages such as hate speech. In this paper, we introduces MemeBLIP2, a light weight multimodal system that detects harmful memes by combining image and text features effectively. We build on previous studies by adding modules that align image and text representations into a shared space and fuse them for better classification. Using BLIP-2 as the core vision-language model, our system is evaluated on the PrideMM datasets. The results show that MemeBLIP2 can capture subtle cues in both modalities, even in cases with ironic or culturally specific content, thereby improving the detection of harmful material.
Keywords Large Language Models (LLMs), Multimodal System, Harmful Content, Artificial Intelligence
@@ 1 Introduction
Detecting harmful memes requires understanding how pictures and text work together. An important early effort was the Hateful Memes Challenge by Kiela et al. [1], which provided a dataset specifically designed to test multimodal hate detection. This dataset was challenging because the text or image alone might seem harmless, but combined they became harmful. This highlighted that analyzing images and text separately is insufficient.
Following this challenge, researchers began using powerful computer models trained on large amounts of visual and textual data. For example, the MOMENTA system [2] not only identified harmful memes but also determined the target of the hate speech by leveraging pre-trained model features. Later, significant improvements came with the CLIP model [11], a system known for effectively matching images with text by learning from vast internet data. Systems like HateCLIPper [3], built specifically around CLIP, showed notable improvements in detecting hateful memes. Recently, MemeCLIP [4] achieved state-of-the-art results by efficiently combining visual and textual information from CLIP, yet subtle meanings and cultural references still posed challenges.
To address these ongoing difficulties, we propose MemeBLIP2—a new, lightweight multimodal system. MemeBLIP2 takes inspiration from MemeCLIP but introduces BLIP-2 [6], a more advanced vision-language model. BLIP-2 better captures subtle details and deeper connections between text and images due to its refined training approach. We expect MemeBLIP2 to significantly improve the detection of hidden or nuanced harmful messages in memes, thus making online spaces safer and more inclusive..
@@ 2 Related Work
Then came the Hateful Memes Challenge [1], a pivotal moment in which meme analysis was framed as a multimodal task. This benchmark showed humans outperformed single-modality AI systems by a large margin, motivating the need for joint text-image understanding. Later work harnessed pretrained vision-language models (VLMs) such as CLIP to fill the gap. For instance, MOMENTA [2] combined CLIP embeddings with social context graphs to improve detection accuracy, and HateCLIPper [3] employed adversarial training to alleviate CLIP’s over-reliance on textual cues. These techniques made significant advances, reporting up to 65–75% accuracy on datasets like the Facebook Hateful Memes dataset. But a key limitations remain. The CLIP-based models, designed to maximize global text-image similarity, commonly failed to connect the fine-grained region-wise interactivity properties (e.g., biased symbols in image corners with neutral captions). Performance gaps also occurred in culturally specific contexts, where sarcastic memes that portray minority groups (e.g., LGBTQ + communities) were often misclassified.
@@
Recent work has attempted to address these issues with disentangled feature learning and community-centric datasets. For instance, have decoupled CLIP’s text and image encoders to improve generalization [4], and datasets such as PrideMM concentrated on underrepresented cultural contexts. Our new routines may be more advanced, but we still have core problems. Static CLIP encoders are ill-suited to the fast-paced evolution of meme styles, including AI-generated visuals, or stylized text. Region-aware methods, which are representative of external object detectors (i.e. YOLO), introduce another level of computational complexity, whilst the opaque nature of the decision-making processes — especially in memes containing sarcasm or cultural metaphors — are a limitation on their practical application for human moderators.
@@
BLIP[5] and its successor BLIP2[6] are unified vision-language models based on the multimodal mixture-of-encoder-decoder (MED) architecture, supporting both understanding and generation tasks. A core component in both models is the image-grounded text encoder, which integrates visual information into textual representations via cross-attention layers inserted between the self-attention and feed-forward modules in each transformer block. A task-specific [Encode] token is added to the input text, and its output embedding serves as the joint image-text representation. Trained with an image-text matching (ITM) loss, this encoder enables fine-grained alignment between modalities, facilitating strong performance in tasks such as image-text retrieval, image captioning, and visual question answering[5].
@@
BLIP2[6] builds on this foundation with enhanced efficiency, scalability, and generative capabilities. It introduces parameter-efficient fine-tuning through lightweight adapter modules, significantly reducing the number of trainable parameters without compromising performance. With diverse training objectives and improved architectural designs, BLIP2 achieves better alignment and more fluent, context-relevant generation. Furthermore, its pre-training on a broader and more diverse image-text dataset strengthens its generalization across domains.
@@ 3 Methodology of MemeBLIP2
In this section, we describe our proposed framework, MemeBLIP2, for harmful meme classification. We rely on the BLIP-2 vision-language model [6] to combine both image and text information. Figure 1 illustrates the overall architecture of MemeBLIP2. Below, we provide the detailed mathematical formulation of each component.
@@ 3.1 Dataset
We starts with the careful preparation of the dataset: PrideMM . Firstly, we load all images and their matching textual annotations from the provided dataset files. Every image is resized to a consistent resolution of 224x224 pixels. This resizing helps keep the inputs uniform, ensuring the model trains smoothly and efficiently. Labels indicating whether memes are harmful or benign are directly taken from the dataset annotations.
@@ 3.2 Linear Projection
After preparing the datasets, the next step involves extracting meaningful features from both visual and textual data using the BLIP-2 model (Salesforce/blip2-opt-2.7b) [6]. The BLIP-2 vision encoder produces visual embeddings, which are 1408-dimensional vectors capturing detailed visual characteristics of each meme image by mapping the i the image $\mathbf{I}_{i}$ to a high-dimensional vector $\mathbf{v}_{i}\in\mathbb{R}^{1408}$.
@@
Although $\mathbf{v}_{i}\in\mathbb{R}^{1408}$ and $\mathbf{u}_{i}\in\mathbb{R}^{768}$ are informative, they live in different embedding spaces. Once we have these embeddings, we can transform them into a shared $1024$-dimensional space. To achieve this, we use linear projection layers that reduce the visual embeddings from 1408 dimensions and increase textual embeddings from 768 dimensions to a common space of 1024 dimensions. In our model, let $x\in\mathbb{R}^{d_{\text{in}}}$ be the input embedding vector(from either the vision or language encoder). The first linear projection is given by
@@
Following the projection, an adapter module [8] further refines the projected features with additional nonlinear transformations. Let $x\in\mathbb{R}^{c}$ be the input to the adapter. First, layer normalization is applied:
@@
Next, $\tilde{x}$ is passed through a bottleneck that reduces the dimensionality to $\tilde{c}=\max(16,\lfloor c/r\rfloor)$ with reduction factor $r=1.5$, followed by GELU activation [9]:
@@ 3.5 Multimodal Fusion
This step encodes direct interaction between each dimension of the visual and textual embeddings. A final feed-forward  [10] block is used to further processes $\mathbf{f}_{i}$ and generates $\mathbf{z}_{i}$, known as the fused multimodal embedding:
@@ 3.6 Classification with MLP-Based Classifier
$\tilde{\mathbf{z}}_{i}=\mathrm{LayerNorm}(\mathbf{z}_{i}),\quad\mathbf{h}_{i}=\mathrm{GELU}(\tilde{\mathbf{z}}_{i}),\quad\mathbf{h}_{i}^{\prime}=\mathrm{Dropout}(\mathbf{h}_{i}),\quad\mathrm{logits}_{i}=\mathbf{W}\,\mathbf{h}_{i}^{\prime}+\mathbf{b}.$
@@
Unlike cosine-based methods, which rely on angular similarity, this MLP-based classifier uses learned transformations and normalization to handle scale differences in the embeddings.
@@ 3.7 Training and Optimization
During training, we optimize the classifier parameters using the AdamW optimizer[7]. The update rule is defined as:
@@
Below is the detailed flow of the MemeBLIP2 methodology outlined :
Figure 1: An overview of our proposed framework, MemeBlip2. We use frozen Q former image encoder from Blip2 and text encoders to create representations for each image-text pair. These representations are passed through linear layers to disentangle the modalities in Blip2’s shared embedding space. We implement Feature Adapters with residual connections for each modality to prevent overfitting. We use a cosine classifier to make MemeBlip2 more robust to imbalanced data. We initialize classifier weights by using Semantic-Aware Initialization to further improve performance.
4 Experiment Plan
@@ 4.2 Experimental Setup
Motivation.
The design of our experimental framework is grounded in the hypothesis that pre-trained vision-language models (VLMs), such as BLIP-2, encapsulate rich cross-modal semantics that can be effectively adapted to domain-specific tasks through lightweight architectural modifications. Prior works like MemeCLIP [shah2024memeclip] have demonstrated that preserving frozen pre-trained encoders while fine-tuning minimal projection and adapter modules achieves competitive performance with high efficiency. Our goal is to validate whether similar strategies remain effective when applied to a stronger backbone (BLIP-2 vs. CLIP) on the PrideMM dataset.
Model Configuration.
@@ 4.3 Baselines
We compare our model against several representative baselines across the unimodal and multimodal spectrum:
@@
  • GPT-4 (Text Only): Zero-shot inference using GPT-4 with prompt-based classification.
  • (To be added): BLIP-2 image-only, text-only, and fusion-only (without adapters) variants.
  • (To be added): MemeBLIP with ablated modules (no adapter, no projection, etc.).
@@ 4.4 Main Results
The proposed MemeBLIP-2 model demonstrates superior performance than all unimodal baselines by reaching 74.8% accuracy and achieving 77.8% AUROC and 76.5% F1-score. The proposed method of combining visual and textual information through effective alignment and fusion proves beneficial for handling noisy meme content that contains rich contextual information.
4.5 Dataset
We conduct our experiments on the PrideMM dataset [4], a benchmark curated for evaluating multimodal models in the context of social media memes. PrideMM contains over 7,000 memes collected from platforms such as Reddit and Twitter, with each sample consisting of an image and a textual caption. The dataset is annotated for harmfulness and offensiveness, posing a challenging binary classification task.
Unlike traditional vision-language datasets, PrideMM includes a significant number of sarcastic, ironic, or culturally nuanced memes, which require the model to jointly reason over visual and textual content to make correct predictions. The dataset is split into training, validation, and test sets following the protocol established in [4], ensuring fair comparison with prior work such as MemeCLIP.
@@ 4.6 Ablation Study on PrideMM Dataset
All modules (baseline) 74.82 ± 1.16 77.81 ± 0.96 76.51 ± 0.94
Remove MLP classifier 73.51 ± 0.71 78.09 ± 0.96 74.42 ± 1.01
w/o MLP + Pre-Output 73.16 ± 1.89 79.22 ± 1.96 73.27 ± 2.35
w/o MLP + Pre + Adapter 73.86 ± 1.70 77.92 ± 1.37 74.01 ± 1.58
Remove All Modules 46.84 ± 1.34 49.73 ± 2.44 43.53 ± 3.38
From Table 2, we observe that the full model achieves the best balance across all metrics. Replacing the MLP classifier with a linear layer leads to a noticeable drop in both accuracy and F1 score, suggesting the effectiveness of non-linear transformations in capturing decision boundaries. Interestingly, AUROC improves slightly, which may indicate that the linear classifier produces more confident outputs on a subset of samples.
Further removal of the pre-output layer and adapters causes gradual performance degradation. When all modules are removed, the model collapses to a shallow baseline and fails to learn meaningful multimodal representations, with accuracy dropping to 46.84% and AUROC approaching random guessing (49.73%). This confirms the importance of architectural depth, modular projection, and modality-specific adaptation for robust multimodal classification, even under simple element-wise fusion.
@@ 4.7 Analysis
Our preliminary experiments suggest that unimodal baselines such as BERT and GPT-4 struggle to disambiguate visual-textual irony and sarcasm present in memes. This validates our hypothesis that effective multimodal processing is essential.
While our model does not surpass MemeCLIP [4] in absolute performance, it offers a complementary perspective on multimodal meme classification. Unlike MemeCLIP, which leverages semantically aligned CLIP embeddings pretrained via contrastive learning, our approach builds upon BLIP2 encoders whose image and text features are independently generated and not inherently aligned.
This design introduces additional challenges in feature fusion and semantic integration, but also enables us to investigate the role of intermediate components such as projection layers, adapters, and fusion strategies. Our ablation study highlights how these modules contribute to bridging the semantic gap between modalities—an aspect less explored in prior CLIP-based work.
Furthermore, our architecture is modular and lightweight, making it adaptable to settings where pretrained contrastive models are unavailable or where encoder-specific customization is required. We believe that our findings offer valuable insights into modality alignment under non-contrastive settings and can guide future efforts in designing flexible and interpretable multimodal systems.
Key insights from the ablation experiments include:
@@
  • Projection layers are critical for semantic alignment between modalities. Removing them results in near-random predictions (AUROC drops from 77.81% to 49.73%).
  • Adapters contribute to representation flexibility and fine-tuning stability, though they are not solely responsible for performance gains.
  • The MLP classifier improves accuracy and F1 by modeling non-linear decision boundaries, with the cosine variant helping mitigate class imbalance.
To further dissect these effects, we conduct a comprehensive ablation study by progressively removing key modules: the MLP classifier, pre-output transformation, adapter blocks, and the projection layers. The performance drops are summarized in Table 2 and visualized in Figure 2.
Figure 2: Performance drop from removing individual modules on the PrideMM dataset.
Removing the classifier results in a moderate decrease in accuracy and F1, highlighting the effectiveness of non-linear transformations. Interestingly, the removal of pre-output and adapter layers leads to relatively minor performance drops, suggesting that they play stabilizing but non-essential roles.
In contrast, removing the projection layers causes a dramatic collapse in all metrics. This underscores a fundamental difference between our model and MemeCLIP: whereas CLIP embeddings are pretrained to be semantically aligned, our BLIP2-based features reside in unaligned spaces. Thus, the projection layers in our model serve not just as dimensional compressors but also as semantic mappers for effective multimodal fusion.
These results confirm that while architectural simplifications may retain partial performance, the full combination of MLP classifier, pre-output transformation, adapters, and projection layers yields the most robust and generalizable model for multimodal meme classification.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Detecting harmful memes requires understanding how pictures and text work together. An important early effort was the Hateful Memes Challenge by Kiela et al. Kiela2020 , which provided a dataset specifically designed to test multimodal hate detection. This dataset was challenging because the text or image alone might seem harmless, but combined they became harmful. This highlighted that analyzing images and text separately is insufficient.
Following this challenge, researchers began using powerful computer models trained on large amounts of visual and textual data. For example, the MOMENTA system Pramanick2021 not only identified harmful memes but also determined the target of the hate speech by leveraging pre-trained model features. Later, significant improvements came with the CLIP model Radford2021 , a system known for effectively matching images with text by learning from vast internet data. Systems like HateCLIPper Kumar2022 , built specifically around CLIP, showed notable improvements in detecting hateful memes. Recently, MemeCLIP Shah2024 achieved state-of-the-art results by efficiently combining visual and textual information from CLIP, yet subtle meanings and cultural references still posed challenges.
To address these ongoing difficulties, we propose MemeBLIP2, a new lightweight multimodal system. MemeBLIP2 takes inspiration from MemeCLIP but introduces BLIP-2 Li2023 , a more advanced vision-language model. BLIP-2 better captures subtle details and deeper connections between text and images due to its refined training approach. We expect MemeBLIP2 to significantly improve the detection of hidden or nuanced harmful messages in memes, thus making online spaces safer and more inclusive.
@@ 2 Related Work
Then came the Hateful Memes Challenge Kiela2020 , a pivotal moment in which meme analysis was framed as a multimodal task. This benchmark showed humans outperformed single-modality AI systems by a large margin, motivating the need for joint text-image understanding. Later work harnessed pretrained vision-language models (VLMs) such as CLIP to fill the gap. For instance, MOMENTA Pramanick2021 combined CLIP embeddings with social context graphs to improve detection accuracy, and HateCLIPper Kumar2022 employed adversarial training to alleviate CLIP’s over-reliance on textual cues. These techniques made significant advances, reporting up to 65–75% accuracy on datasets like the Facebook Hateful Memes dataset. But a key limitations remain. The CLIP-based models, designed to maximize global text-image similarity, commonly failed to connect the fine-grained region-wise interactivity properties (e.g., biased symbols in image corners with neutral captions). Performance gaps also occurred in culturally specific contexts, where sarcastic memes that portray minority groups (e.g., LGBTQ + communities) were often misclassified.
@@
Recent work has attempted to address these issues with disentangled feature learning and community-centric datasets. For instance, have decoupled CLIP’s text and image encoders to improve generalization Shah2024 , and datasets such as PrideMM concentrated on underrepresented cultural contexts. Our new routines may be more advanced, but we still have core problems. Static CLIP encoders are ill-suited to the fast-paced evolution of meme styles, including AI-generated visuals, or stylized text. Region-aware methods, which are representative of external object detectors (i.e.YOLO), introduce another level of computational complexity, whilst the opaque nature of the decision-making processes — especially in memes containing sarcasm or cultural metaphors — are a limitation on their practical application for human moderators.
@@
BLIPLi2022 and its successor BLIP2Li2023 are unified vision-language models based on the multimodal mixture-of-encoder-decoder (MED) architecture, supporting both understanding and generation tasks. A core component in both models is the image-grounded text encoder, which integrates visual information into textual representations via cross-attention layers inserted between the self-attention and feed-forward modules in each transformer block. A task-specific [Encode] token is added to the input text, and its output embedding serves as the joint image-text representation. Trained with an image-text matching (ITM) loss, this encoder enables fine-grained alignment between modalities, facilitating strong performance in tasks such as image-text retrieval, image captioning, and visual question answeringLi2022 .
@@
BLIP2Li2023 builds on this foundation with enhanced efficiency, scalability, and generative capabilities. It introduces parameter-efficient fine-tuning through lightweight adapter modules, significantly reducing the number of trainable parameters without compromising performance. With diverse training objectives and improved architectural designs, BLIP2 achieves better alignment and more fluent, context-relevant generation. Furthermore, its pre-training on a broader and more diverse image-text dataset strengthens its generalization across domains.
@@ 3 Methodology of MemeBLIP2
In this section, we describe our proposed framework, MemeBLIP2, for harmful meme classification. We provide the detailed mathematical formulation of each component. The detailed flow of the MemeBLIP2 methodology is outlined in Figure  1.
Figure 1: An overview of our proposed framework, MemeBlip2. We use frozen Q former image encoder and text encoders from BLIP-2 to create representations for each image-text pair. These representations are passed through linear layers to disentangle the modalities in BLIP-2’s shared embedding space. We implement Feature Adapters with residual connections for each modality to prevent overfitting. We use a MLP-based classifier to make MemeBLIP2 more robust to imbalanced data. We initialize MLP-based classifier weights by using Semantic-Aware Initialization to further improve performance.
@@ 3.1 Dataset
We conduct our experiments on the PrideMM dataset Shah2024 , a benchmark curated for evaluating multimodal models in the context of social media memes. PrideMM contains over 7,000 memes collected from platforms such as Reddit and Twitter, with each sample consisting of an image and a textual caption. The dataset is annotated for harmfulness and offensiveness, posing a challenging binary classification task.
Unlike traditional vision-language datasets, PrideMM includes a significant number of sarcastic, ironic, or culturally nuanced memes, which require the model to jointly reason over visual and textual content to make correct predictions. The dataset is split into training, validation, and test sets following the protocol established in Shah2024 , ensuring fair comparison with prior work such as MemeCLIP.
We start with the careful preparation of the dataset by loading all images and their matching textual annotations from the provided dataset files. Every image is resized to a consistent resolution of 224x224 pixels. This resizing helps keep the inputs uniform, ensuring the model trains smoothly and efficiently. Labels indicating whether memes are harmful or benign are directly taken from the dataset annotations.
@@ 3.2 Linear Projection
After preparing the datasets, the next step involves extracting meaningful features from both visual and textual data using the BLIP-2 model (Salesforce/blip2-opt-2.7b) Li2023 . The BLIP-2 vision encoder produces visual embeddings, which are 1408-dimensional vectors capturing detailed visual characteristics of each meme image by mapping the i the image $\mathbf{I}_{i}$ to a high-dimensional vector $\mathbf{v}_{i}\in\mathbb{R}^{1408}$.
@@
Although $\mathbf{v}_{i}\in\mathbb{R}^{1408}$ and $\mathbf{u}_{i}\in\mathbb{R}^{768}$ are informative, they live in different embedding spaces. Once we have these embeddings, we can transform them into a shared 102410241024-dimensional space. To achieve this, we use linear projection layers that reduce the visual embeddings from 1408 dimensions and increase textual embeddings from 768 dimensions to a common space of 1024 dimensions. In our model, let $x\in\mathbb{R}^{d_{\text{in}}}$ be the input embedding vector(from either the vision or language encoder). The first linear projection is given by
@@
Following the projection, an adapter module Houlsby2019 further refines the projected features with additional nonlinear transformations. Let $x\in\mathbb{R}^{c}$ be the input to the adapter. First, layer normalization is applied:
@@
Next, $\tilde{x}$ is passed through a bottleneck that reduces the dimensionality to $\tilde{c}=\max(16,\lfloor c/r\rfloor)$ with reduction factor $r=1.5$, followed by GELU activation Hendrycks2016 :
@@ 3.5 Multimodal Fusion
This step encodes direct interaction between each dimension of the visual and textual embeddings. A final feed-forward  Vaswani2017 block is used to further processes $\mathbf{f}_{i}$ and generates $\mathbf{z}_{i}$, known as the fused multimodal embedding:
@@ 3.6 Classification with MLP-Based Classifier
$\tilde{\mathbf{z}}_{i}$ $=\mathrm{LayerNorm}(\mathbf{z}_{i}),\quad\mathbf{h}_{i}=\mathrm{GELU}(\tilde{\mathbf{z}}_{i}),$
$\mathbf{h}_{i}^{\prime}$ $=\mathrm{Dropout}(\mathbf{h}_{i}),\quad\mathrm{logits}_{i}=\mathbf{W}\,\mathbf{h}_{i}^{\prime}+\mathbf{b}.$
@@
Unlike cosine-based methods, which rely on angular similarity, this MLP-based classifier uses learned transformations and normalization to handle scale differences in the embeddings. Layer normalization stabilizes training, and the GELU nonlinearity plus dropout relieve the problem of overfitting. As our empirical results show, these added nonlinear transformations allow the MLP to model subtle semantic cues more effectively, especially in imbalanced datasets, thus outperforming the cosine classifier in our experiments.
@@ 3.7 Training and Optimization
During training, we optimize the classifier parameters using the AdamW optimizerLoshchilov2018 . The update rule is defined as:
@@
4 Experiments
@@ 4.2 Experimental Setup
Motivation
The design of our experimental framework is grounded in the hypothesis that pre-trained vision-language models (VLMs), such as BLIP-2, encapsulate rich cross-modal semantics that can be effectively adapted to domain-specific tasks through lightweight architectural modifications. Prior works like MemeCLIP Shah2024 have demonstrated that preserving frozen pre-trained encoders while fine-tuning minimal projection and adapter modules achieves competitive performance with high efficiency. Our goal is to validate whether similar strategies remain effective when applied to a stronger backbone (BLIP-2 vs. CLIP) on the PrideMM dataset.
Model Configuration
@@ 4.3 Baselines
We compare our model MemeBLIP2 against several representative baselines across the unimodal and multimodal spectrum as well as the orginal MemeCLIP:
@@
  • GPT-4 (Text Only): Zero-shot inference using GPT-4o with prompt-based classification.
  • GPT-4 (Text + Image): GPT-4o with multi model input enabled.
  • MemeCLIP: Trained and tested using the same settings as the original paper.
@@ 4.4 Main Results
The proposed MemeBLIP-2 model demonstrates superior performance than all unimodal baselines by reaching 77.5% accuracy and achieving 81.8% AUROC and 79.0% F1-score. It also outperforms MemeCLIP, the paper that we mainly refer, by 3.1% in accuracy and 6.0% in F1-score, but the AUROC is slightly lower by 1.6% .
In a word, the proposed method of combining visual and textual information through effective alignment and fusion proves beneficial for handling noisy meme content that contains rich contextual information.
4.5 Gradient Diagnosis
Figure 2: Gradient standard deviation of individual parameters across training epochs. Notable fluctuations are observed in the classifier’s final weight and bias terms.
To better understand the training dynamics of our model, we monitored the gradient standard deviation (grad_std) of each parameter throughout training. Figure 2 shows that while most modules—such as the adapters and projection layers—maintain relatively stable gradient distributions, the classifier head exhibits significant variance. In particular, the classifier.7.weight and classifier.7.bias parameters display repeated spikes, exceeding 10,000 in gradient standard deviation during multiple epochs.
This behavior indicates heightened sensitivity or instability in the classifier head, which aligns with our earlier ablation findings: removing or simplifying the classifier leads to notable performance shifts. These fluctuations suggest that further regularization, layer-specific learning rates, or gradient clipping might be beneficial, especially for deeper or more nonlinear components such as MLP classifiers.
Overall, this analysis complements our architectural ablations by highlighting potential optimization bottlenecks and emphasizing the importance of balanced gradient flow across modular components.
@@ 4.6 Ablation Study on PrideMM Dataset
All modules (baseline) 76.90 ± 0.55 80.80 ± 0.96 78.39 ± 0.61
Remove MLP classifier 76.90 ± 0.21 80.16 ± 0.81 78.67 ± 0.55
w/o MLP + Pre-Output 76.46 ± 0.41 79.28 ± 0.31 77.71 ± 0.82
w/o MLP + Pre + Adapter 75.44 ± 0.36 76.79 ± 3.90 77.25 ± 1.08
Remove All Modules 49.56 ± 4.07 48.22 ± 1.73 46.42 ± 9.14
From Table 2, we observe that the full model achieves the best AUROC and is among the best in accuracy and F1 score, indicating a strong balance across evaluation metrics. Removing the MLP classifier has minimal effect on accuracy, but slightly decreases AUROC, suggesting that the classifier architecture helps produce more calibrated predictions.
Further removing the pre-output layer causes a moderate drop across all metrics, particularly AUROC and F1. This implies the pre-output transformation contributes meaningfully to downstream decision quality. When adapters are also removed, the degradation becomes more evident, supporting their role in enhancing modality-specific representation learning. Finally, removing all major modules—including the projection layers—leads to a dramatic performance collapse, with accuracy dropping to 49.56% and AUROC near random (48.22%). This underscores the necessity of layered projections and modular components for effective multimodal integration.
@@ 4.7 Analysis
Our experiments confirm that unimodal baselines such as BERT and GPT-4 struggle with interpreting visually grounded irony and sarcasm in memes. This validates our hypothesis that effective multimodal integration is essential for nuanced meme classification.
While our model does not surpass MemeCLIP Shah2024 in absolute performance, it provides an alternative architecture for multimodal fusion. Unlike MemeCLIP—which uses contrastively pre-aligned CLIP embeddings—our method builds on BLIP2, where vision and language features are extracted independently and not inherently aligned. This introduces fusion challenges, but also offers flexibility in modular design and encoder selection.
The ablation study (Table 2) reveals several insights. When all modules are intact, the model achieves the highest AUROC (80.80%) and maintains strong accuracy (76.90%) and F1 (78.39%). Removing only the MLP classifier has little impact on accuracy, but leads to a mild AUROC decline—indicating its role in shaping output calibration rather than feature quality. Further removal of the pre-output transformation causes additional degradation across metrics, particularly F1, showing its contribution to the final representation.
Removing adapter modules causes a larger performance drop, suggesting they play a meaningful role in modality-specific transformation. Finally, removing all major modules—including projection layers—leads to a performance collapse, confirming their critical role in aligning unaligned modalities. AUROC drops from 80.80% to 48.22%, close to random, and F1 drops by over 30%.
Key insights from the updated ablation experiments:
@@
  • Projection layers are indispensable for cross-modal alignment. Removing them causes near-random predictions (AUROC: 80.80% $\rightarrow$ 48.22%).
  • Adapters help improve feature expressiveness and adaptation, contributing to stable and generalizable learning.
  • The MLP classifier improves calibration and downstream discrimination, although its removal does not drastically harm performance.
  • The pre-output transformation provides additional abstraction and improves F1, supporting its inclusion despite moderate standalone impact.
To better understand each component’s contribution, we visualize the performance degradation across accuracy, AUROC, and F1 score in Figure 3. This incremental analysis confirms that while some modules (e.g., classifier, pre-output) offer moderate benefits, the adapter and projection layers play a pivotal role in preserving cross-modal semantics and stable training dynamics.
Figure 3: Incremental performance drop across metrics (ACC, AUROC, F1) when removing each module. Projection layers have the most significant impact, highlighting their role in modality alignment.
Taken together, our findings emphasize the importance of intermediate layers in bridging the semantic gap between vision and language features under non-contrastive, independently encoded settings. These results suggest that while partial simplifications may preserve some performance, the combination of modular design, MLP-based classification, and modality-specific transformations produces the most robust architecture for multimodal meme classification.